OpenStudy (anonymous):

find u and v where sinu=2sinv and cosu+2cosv=2

OpenStudy (asnaseer):

try squaring the first one and then using the rule:\[\sin^2(x)=1-\cos^2(x)\]

OpenStudy (anonymous):

sure, just managed it with cos2x=1-sin^2x but it felt a bit over the top; thought there might be a better way

OpenStudy (asnaseer):

I think thats the right approach.

OpenStudy (anonymous):

coool

OpenStudy (anonymous):

thanks

OpenStudy (asnaseer):

yw